Complete the following calculations in an Excel file. Be sure to show and label your work for full credit.

Grizzly Company produces stuffed bears. Based on its history, it typically uses 3 pounds of material for each bear, costing $6 per pound. Each bear is expected to have direct labor of 1 hour at a rate of $10.50. In the month of January, the company produced 155,000 bears. They had direct materials of 470,000 pounds, costing on average $5.95. Direct labor was 143,000 hours at an average wage of $11.00.

  1. Compute the materials price variance, materials usage variance, and the total materials variance.
  2. Compute the labor rate variance, labor efficiency variance, and the total labor variance.
  3. Assume the company only investigates variances if they are $7,500 or more. Should it investigate? And if so, would the answer be different if it were to look only at the total variances?
Answer #1
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($6/Pound - $5.95/Pound ) × 470000 Pound = $23500 F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((3Pound X 155000 Unit)-470000 Pound )X $6 = $30000 UF
Total Material Variance= MPV+MQV
=$23500 F+$30000UF=$6500 UF
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($10.50/hour - $11/hour) × 143000 Hours= $71500 UF
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard Rate per Hour
= (1 Hour × 155000 Unit - 143000) × $10.50 = $126000 F
Total Labour Variance= LRV+LEV
=$71500U F+$126000F=$54500 F

Yes company should investigate because all the variance is more than $7500.

Yes, because, it will be focus on total Variance, Material variance will not be investigated due to lesser than $7500